OpenStudy (anonymous):

12. which equation is an identity? A. 7 - (9x + 3)= -9x - 4 B. 6m - 5 = 7m + 5 - m C. 10p + 6 - p= 12p - 3(p-2) D. 3y + 2=3y - 2 13. which equation has no solution? A. 5v + 3 = 5v - 5 B. 2x - 4 = 2x + 7 - x C. 2y + 5 - y= 2y - 3(y-6) D. 7z + 9 = -7z -6

OpenStudy (lieutenantgeneral):

12.) c. 13.) a.

OpenStudy (mathstudent55):

Solve each equation. If the variables cancel out, and you get a true statement, it is an identity. If the variables cancel out, and you get a false statement, it has no solution.

OpenStudy (mathstudent55):

Let's do choice A in problem 12 together. 7 - (9x + 3)= -9x - 4 Distribute the negative sign on the left side: 7 - 9x - 3 = -9x - 4 Combine like terms: 4 - 9x = -9x - 4 Add 9x to both sides: 4 = -4 The variables canceled out, and we got a false statement. That means choice A has no solution.